What are the values of these sums, where S={1,3,5,7}?
(a)∑(jϵS) j
(b)∑(jϵS) j^2
(c)∑(jϵS) (1/j)
(d)∑(jϵS) 1
1
Expert's answer
2020-10-14T16:56:10-0400

Sum a.)

"1+3+5+7=16"

Sum b.)

"1^2+3^2+5^2+7^2=1+9+25+49=84"

Sum c.)

"1\/1+1\/3+1\/5+1\/7="

"105\/105+35\/105+21\/105+15\/105="

"176\/105=1.6762"

Sum d.)

"1+1+1+1=4"